OpenStudy (ineedhelp10):

An adult pass for a county fair costs $2 more than a children’s pass. When 378 adult and 214 children’s passes were sold, the total revenue was $2,348. Find the cost of an adult pass.

OpenStudy (ineedhelp10):

@Directrix help please

OpenStudy (ineedhelp10):

@Compassionate if you're not busy can you please help

OpenStudy (godlovesme):

Adult pass = 2+c 368a+214c=2348 That's the info we have so far but idk how to set it up :/

OpenStudy (godlovesme):

Do u have options @ineedhelp10

OpenStudy (ineedhelp10):

@Directrix do you have an idea on how to set it up? and thanks @Godlovesme

OpenStudy (godlovesme):

Np but do u have choices?

OpenStudy (ineedhelp10):

no its not multiple choice

Directrix (directrix):

378 Adult Tickets Sold 214 Child Tickets Sold Income from Adult Tickets: 378 * A Income from Child Tickets: 214 * C But, a child's ticket costs 2 dollars less than an adult ticket so C = (A - 2) 378 * A + 214* (A - 2) = 2348 ---> Equation, solve it for A. @ineedhelp10

OpenStudy (godlovesme):

http://www.algebra.com/algebra/homework/word/mixtures/Mixture_Word_Problems.faq.question.193138.html Try this :) I really really hope helps

OpenStudy (ineedhelp10):

okay I'll try that website out @Godlovesme

OpenStudy (godlovesme):

this helps*** @ineedhelp10

OpenStudy (godlovesme):

K :)

OpenStudy (ineedhelp10):

so i just have to solve that equation @Directrix

OpenStudy (ineedhelp10):

thnx ;) @Godlovesme

OpenStudy (godlovesme):

Anytime ^_^ @ineedhelp10

OpenStudy (godlovesme):

@ineedhelp10 could u close this question to let every1 kno that u don't need help :)